Find the integral surface of the equation 4yzp+q+2y=0 passing through

y2+z2=2,x+z=1


1
Expert's answer
2021-11-01T13:16:00-0400

"\\\text{Given:}\ 4 y z p+q+2y=0 \ldots (1)
\\\text{Given curve is}\ y^{2}+z^{2}=2, x+z=1 \ldots (2)
\\\text{The Lagrange’s auxiliary equations of (1) are}
\\\frac{d x}{4 y z}=\frac{d y}{1}=\frac{d z}{-2 y} \quad \ldots (3)
\\\text{Taking the first and third fraction of (3), we have}
\\d x+2 z d z=0, \text{so that}\ x+z^2=c_{1} \cdots (4)
\\\text{Taking the last two fractions of (3), we have}
\\d z+2 y d y=0 \text { so that }\ z+y^2=c_{2} \ldots (5)
\\\text{Adding (4) and (5),} 
\\ x+z^2+ z+y^2= c_{1}+ c_{2}
\\\Rightarrow y^{2}+z^{2}+x+z=c_{1}+c_{2}
\\\Rightarrow 2+1=c_{1}+c_{2}
\\\Rightarrow c_{1}+c_{2} =3 
\\\text{The equation of required surface is}
\\x+z^2+ z+y^2 =3
\\\Rightarrow y^{2}+z^{2}+x+z-3=0"